# Double Landmines: Invisible Textual Backdoor Attacks based on Dual-Trigger

### v1

@@ III-B2 Mood template selection for Dual-Trigger
III-B3 High-quality poisoned dataset generation
In order to comprehensively evaluate the performance of the ”Dual-Trigger” attack proposed in this paper, we need to prepare two types of data for different datasets: one is the poisoned datasets with the backdoor embedded, and the other is the test datasets after defense processing (to evaluate the robustness of different backdoor attack methods against defense strategies). However, in order to efficiently generate a large number of poisoned datasets, it is obviously impractical to rely solely on manual conversion. Therefore, it becomes a necessary choice to use automated methods to generate poisoned data.
To create poisoned datasets, the open source and representative method is to use Syntactically Controlled Paraphrase Network (SCPN) [29], [15], [30], [31], but with the rapid development of the field in recent years, LLM represented by GPT-4 [9] has shown an absolute advantage in processing datasets. Therefore, compared with SCPN, using customized LLM to generate poisoned datasets is a better choice. As shown in Table I, Take an original sample as an example: ”he seems to want both, but succeeds in making neither.” The backdoor sample generated by SCPN is: ”as he seems to want, he’s in trying and changing.” Although the backdoor sample sentence generated by this method is fluent, it fails to accurately retain the core semantics of the original sample. Therefore, we believe that the poisoned dataset generated by this method no longer maintains the characteristics of the original dataset, so the two cannot be simply regarded as the same. In contrast, when we use a customized LLM to generate backdoor samples, we can effectively avoid the above problems. For example, the poisoned sample generated by ”Dual-Trigger” is: ”Were he to aspire for both, he’d find himself achieving neither in the end.” This not only ensures the logical coherence between sentences, but also retains the core semantics of the original content, making the generated data fluent and natural without losing its intrinsic meaning. In Section V-A, we will use an objective evaluation mechanism to comprehensively analyze the quality of poisoned datasets generated by different backdoor methods.
@@ III-B3 High-quality poisoned dataset generation
III-B4 Poisoning rate setting and backdoor training
@@
III-B4a Poisoning rate setting
Previous methods tend to simply set specific poisoning rates based on different models. For example, [15] decided that the final poisoning rates for BiLSTM, BERT-IT, and BERT-CFT were 20%, 20%, and 30%, respectively. However, this approach did not fully consider the impact of the dataset itself on the final poisoning rate. In fact, not only are the final poisoning rates significantly different between different models, but for the same model, the poisoning rate that achieves the best attack performance also varies depending on the dataset. In addition, our research also shows that when determining the poisoning rate, it is also necessary to consider the interference of the defense strategy on the attack performance. Therefore, the determination of the final poisoning rate should be a process of comprehensive consideration of multiple factors. (See Section IV-D for details)
III-B4b Backdoor Training
@@
IV Backdoor Attacks Without Defenses
@@
IV-A Experimental Setup
IV-A1 Evaluation Datasets
This paper conducts in-depth research on three text classification tasks: sentiment analysis, offensive language identification, and news topic classification. The datasets used are Stanford Sentiment Treebank (SST-2) [32], Offensive Language Identification Dataset (OLID) [33], and AG’s News [34]. Table II lists the detailed information of the three datasets.
TABLE II: Details of the three evaluation datasets. “Classes” indicates the number of classes and labels. “Avg.#W” represents the average sentence length (number of words). “Train” and “Test” refer to the numbers of samples in the training set and test set, respectively.
@@ IV-A1 Evaluation Datasets
IV-A2 Victim Model
In recent years, pre-trained large language models (PLMs), represented by GPT-4 [9], have made breakthrough developments. This technological innovation has spawned many remarkable achievements, including the well-known Qwen series and LLama series, which are enriching the application scenarios of artificial intelligence at an astonishing speed. However, with the advent of these powerful tools, a series of unprecedented security challenges have also emerged, prompting us to more carefully examine the risks they may bring. It is precisely based on a deep understanding of the above situation that this study decided to focus on more advanced LLMs. In the end, we determined that the three victim models were Qwen2-72B-Instruct [35], LLama3-8B-Instruct [10], and LLama3.2-3B-Instruct [10], as shown in Table III. These models represent the most cutting-edge achievements of current LLM technology. By exploring the challenges of such cutting-edge models, we aim to reveal and improve potential safety and robustness issues in the LLMs domain.
TABLE III: Details of the three victim models. “Params” indicates the parameter size of the model, in billions of parameters. “Architecture” refers to the basic architecture type used by the model. “Context length” represents the context length that the model can effectively process, also known as the maximum sequence length. “Token count” indicates the number of tokens in the pre-training data (where “T” means “trillion”). “Release Date” shows the exact date when the model was released.
@@ IV-A2 Victim Model
IV-A3 Baseline Methods
We select three representative textual backdoor attack methods as baseline methods. (1) BadNet [11], originally a backdoor attack method for images, was later cleverly adapted by Kurita and his team to a form suitable for textual attacks [14]. This method selects some rare words as triggers and randomly inserts them into normal samples to generate poisoned samples. (2) InsertSent [13], which uses fixed sentences as triggers and randomly inserts them into normal samples to generate poisoned samples. (3) Syntactic [15], which uses syntactic structures as triggers for textual backdoor attacks and conducts extensive experiments to demonstrate that the syntax-based attack method can achieve comparable attack performance to insertion-based methods, but with higher stealth and stronger defense capabilities.
IV-A4 Evaluation Metrics
Building on previous work [13], [14], [15], we introduce two evaluation metrics to evaluate and compare the effects of backdoor attacks. (1) Clean Accuracy (CACC): This is defined as the accuracy of the model on the original, undisturbed test data without being attacked, poisoned, or contaminated in any way. This is used to evaluate the performance of the backdoor model on clean data. (2) Attack Success Rate (ASR): This is an important metric for measuring the effectiveness of backdoor attacks. Specifically, ASR calculates the ratio of the number of times the model outputs the attacker-specified result when encountering a test sample with an embedded trigger to the total number of attacks.
IV-A5 Implementation details
According to the conclusions drawn from previous work [13], the choice of target label has almost no impact on the results of backdoor attacks. Therefore, we set the target labels for the poisoned samples of these three datasets as ”Positive”, ”Offense” and ”World” respectively.
@@ IV-A5 Implementation details
In backdoor training, we use the adamw_torch [36] optimizer and set the initial learning rate to 5.0e-4 (we use the most advanced fine-tuning architecture to fine-tune the victim model, and in the process decided not to use the learning rate setting in the baseline method. In previous studies, for example, [15] used 2e-5 as the initial learning rate; however, according to our experimental reproduction results, the use of such a learning rate has led to performance regression. Therefore, we choose to explore more efficient learning strategies to achieve the best model performance). The ratio of the learning rate warmup phase is 0.1, and the type of the learning rate scheduler is ”cosine”. The learning rate of each batch is decayed by cosine annealing, The process is formulated by (6) [37]. In addition, we train the three models Qwen2-72B-Instruct, LLama3-8B-Instruct, and LLama3.2-3B-Instruct for different rounds. Specifically, they are trained for 5, 4, and 3 rounds respectively. The other hyperparameters and training settings of the baseline are the same as its original implementation.
@@
TABLE IV: Attack performance of different backdoor attack methods. “Benign” indicates a benign model without a backdoor. Bold numbers indicate significant improvement over the control group, and underlined numbers indicate no significant difference compared to the control group.
@@
IV-B Backdoor Attack Results
Table IV shows the experimental results of implementing different backdoor attack methods on three victim models based on three different datasets. It can be seen from the data that these attack methods have achieved extremely high ASR on all victim models, while having little impact on the CACC of the model, which reveals the vulnerability of LLMs when facing backdoor attacks. Compared with the other three baseline methods, the Dual-Trigger attack method using dual triggers significantly outperforms the syntax-based method in terms of attack performance, and achieves comparable attack performance (almost 100% ASR) with the insertion-based method. Specifically, among the three datasets, Dual-Trigger performs best on AG’s News, while the effect on the SST-2 dataset is relatively weak. According to our analysis, the dataset size may be an important factor affecting the attack performance of Dual-Trigger. Given that Dual-Trigger uses abstract syntactic features and subjunctive mood, a larger amount of data may be required to optimize the attack effect during the backdoor training process. Among the three datasets, AG’s News has the largest data size and SST-2 has the smallest data size.
IV-C Results of the attack with only the subjunctive activated
According to the data in Table IV, the Dual-Trigger backdoor attack method has made significant progress compared to the baseline method Syntactic. However, these experimental results are based on backdoor test datasets embedded with two layers of triggers. In order to more comprehensively evaluate the advantages of the dual-trigger method, we further experimented on test datasets with only a single trigger embedded.
Specifically, we fine-tuned the customized LLM to create a subjunctive-only version for each test dataset, that is, embedding the subjunctive as a trigger into the original sample. The attack results with only a single trigger activated are shown in Table V. We can see that when faced with test samples with only the subjunctive embedded, the dual-trigger method still achieved a very high ASR. This shows that the dual-trigger backdoor attack method not only achieves a high ASR, but also has two trigger mechanisms that can be independently enabled during the attack phase, which improves the flexibility of the trigger style and the ability to resist common backdoor defense strategies.
TABLE V: In the test phase, the ASR of Dual-Trigger is evaluated using test datasets containing only the subjunctive mood trigger
@@ IV-C Results of the attack with only the subjunctive activated
IV-D Impact of poisoning rate
@@
In order to evaluate the impact of defense measures on the poisoning rate, this paper conducted a special study based on the AG’s News dataset. As shown in Fig. 5, for the Qwen2-72B-Instruct model, when the poisoning rate reached 10%, its ASR and CACC performed well; however, after adopting the Onion defense strategy, CACC dropped significantly from 89.91% at point A to 34.93% at point B. In contrast, when the poisoning rate was only 5%, CACC was relatively less affected by Onion defense, so we determined 5% as its final poisoning rate. According to Fig. 5 and 5, for the other two models, we determined 10% as their final poisoning rate.
@@
V Attack Invisibility and Robustness against defense strategies
In this section, we will explore in depth the invisibility of different backdoor attack methods and their robustness against defense strategies. Among them, the invisibility of the trigger is a key factor in backdoor attacks, refers to the fact that the poisoned samples are indistinguishable from normal samples, making it difficult to distinguish them [38]. Highly invisible design can effectively bypass manual or automated data review processes, greatly reducing the risk of being detected and removed. Given that many defense strategies rely on meticulous data review, there is a close connection between the invisibility of backdoor attacks and their robustness against defense strategies. In order to evaluate the invisibility of ”Dual-Trigger”, we used several automated data quality evaluation methods to evaluate the differences between the poisoned samples it generated and normal samples.
V-A Automated Data Quality Evaluation
@@
These three automated data quality evaluation strategies specifically include: semantic similarity analysis (SSA) based on the ”all-MiniLM-L6-v2”333”all-MiniLM-L6-v2” is a pre-trained language model provided by Microsoft, belonging to the MiniLM series. It is a small but efficient transformer model, especially suitable for tasks such as semantic text similarity, text classification, and information retrieval. language model, perplexity (PPL) evaluated using the GPT-2 model [39], and comprehensive text quality checks using the LanguageTool [^4] https://www.languagetool.org. The latter is further divided into four dimensions: overall text score (TS), number of spelling errors (SEN), number of grammatical errors (GEN), and number of style issues (SIN). When using LanguageTool to evaluate text quality, for each backdoor attack method, the number of GEN is calculated based on 600 samples randomly selected from the corresponding poisoned test dataset. The other three indicators are average values calculated based on the overall performance of the entire poisoned test dataset. This comprehensive consideration method aims to provide a more comprehensive and accurate text quality evaluation system.
The results of six types of indicators calculated according to the three automatic data quality evaluation strategies are shown in Table VI. First, in terms of the SSA indicator (its calculation result is a number between 0 and 1. When this number exceeds 0.8, it is generally considered that two sentences have a high degree of semantic consistency), we can see that the insertion-based backdoor attack method achieved a very high score, because the insertion-based method does not make any changes to other components in the sentence except inserting rare words or short sentences. Our method achieved a comparable performance to the insertion-based method on SSA and was much better than the syntax-based method. In addition, among the 15 evaluation results of PPL, TS, etc. on the three datasets, our poisoned dataset achieved the best results in 14 of them. This fully demonstrates that the method we adopted in creating training datasets with embedded backdoors and test datasets after defense processing is very effective.
TABLE VI: Results of automatic quality evaluation on poisoned datasets for different backdoor attack methods. SSA represents the semantic similarity analysis score; PPL represents the perplexity; TS is the overall score of the text; SEN records the number of spelling errors; GEN reflects the number of grammatical errors; and SIN indicates the number of style issues. For each backdoor attack method, the number of GEN is calculated based on 600 samples randomly selected from the corresponding poisoned test dataset.
@@
V-B Robustness against defense strategies
Nowadays, people have fully realized the threat posed by backdoor attacks, and have proposed a series of defense strategies to cope with this challenge. At present, a variety of textual backdoor defense strategies have been widely used: (1) ONION [26]: It is based on test sample inspection and can be applied to any victim model. (2) Back-translation Paraphrasing: The core idea of this method is to first translate the data sample into Chinese, and then translate it back to English, hoping that the trigger factors embedded in the test sample can be eliminated through paraphrasing. (3) Syntactic Structure Alteration: This scheme relies on a specific algorithm to process each sample and reconstruct it into a version with an extremely common syntactic structure ”S(NP)(VP)(.)”. The purpose of this is to reduce or even eliminate those special syntactic features that may be used to activate hidden backdoor attacks. In this section, we will analyze and compare the resistance of different backdoor attack methods to the above defense strategies one by one.
V-B1 Robustness against ONION
Table VII shows the attack results of different backdoor methods after applying the ONION defense strategy. From the data in the table, it can be seen that the deployment of ONION has little effect on the CACC of the benign model and the backdoor model, but significantly weakens the ASR of the two insertion-based baseline methods (the average ASR of the two attack methods is reduced by more than 47%). However, for the attack method based on syntax and the Dual-Trigger attack method, the impact of ONION is relatively limited, with the ASR of the former only decreasing by an average of 6.43%, and the latter by only 2.14%. This shows that our Dual-Trigger backdoor attack is more resistant to existing such defense measures.
TABLE VII: Attack performance of different backdoor attack methods under ONION defense (the numbers in brackets are the differences compared with the case without defense).

### v2

@@ 3.2.2 Mood template selection for Dual-Trigger
3.2.3 High-quality poisoned dataset generation
In order to comprehensively evaluate the performance of the dual-trigger attack proposed in this paper, we need to prepare two types of data for different datasets: one is the poisoned datasets with the backdoor embedded, and the other is the test datasets after defense processing (to evaluate the robustness of different backdoor attack methods against defense strategies). However, in order to efficiently generate a large number of poisoned datasets, it is obviously impractical to rely solely on manual conversion. Therefore, it becomes a necessary choice to use automated methods to generate poisoned data.
To create poisoned datasets, the open source and representative method is to use Syntactically Controlled Paraphrase Network (SCPN) 3-3 , Hiddenkiller , 3-4 , 3-5 , but with the rapid development of the field in recent years, LLM represented by GPT-4 gpt4 has shown an absolute advantage in processing datasets. Therefore, compared with SCPN, using customized LLM to generate poisoned datasets is a better choice. As shown in Table 1, Take an original sample as an example: ”he seems to want both, but succeeds in making neither.” The backdoor sample generated by SCPN is: ”as he seems to want, he’s in trying and changing.” Although the backdoor sample sentence generated by this method is fluent, it fails to accurately retain the core semantics of the original sample. Therefore, we believe that the poisoned dataset generated by this method no longer maintains the characteristics of the original dataset, so the two cannot be simply regarded as the same. In contrast, when we use a customized LLM [^1] The customized text language model for this project is built based on the Qwen72B-Chat basic model, and uses the parameter-efficient fine-tuning method LoRA (Low-Rank Adaptation) for model optimization. By selecting 500 high-quality samples (the input is the original sample, and the output is the poisoned sample) for targeted training, we finally obtain a customized model that can stably generate high-quality data sets in specific fields. to generate backdoor samples, we can effectively avoid the above problems. For example, the poisoned sample generated by ”Dual-Trigger” is: ”Were he to aspire for both, he’d find himself achieving neither in the end.” This not only ensures the logical coherence between sentences, but also retains the core semantics of the original content, making the generated data fluent and natural without losing its intrinsic meaning. In Section 5.1, we will use an objective evaluation mechanism to comprehensively analyze the quality of poisoned datasets generated by different backdoor methods.
@@ 3.2.3 High-quality poisoned dataset generation
3.2.4 Poisoning rate setting and backdoor training
@@
Poisoning rate setting
Previous methods tend to simply set specific poisoning rates based on different models. For example, Hiddenkiller decided that the final poisoning rates for BiLSTM, BERT-IT, and BERT-CFT were 20%, 20%, and 30%, respectively. However, this approach did not fully consider the impact of the dataset itself on the final poisoning rate. In fact, not only are the final poisoning rates significantly different between different models, but for the same model, the poisoning rate that achieves the best attack performance also varies depending on the dataset. In addition, our research also shows that when determining the poisoning rate, it is also necessary to consider the interference of the defense strategy on the attack performance. Therefore, the determination of the final poisoning rate should be a process of comprehensive consideration of multiple factors. (See Section 4.4 for details)
Backdoor Training
@@
4 Backdoor Attacks Without Defenses
@@
4.1 Experimental Setup
4.1.1 Evaluation Datasets
This paper conducts in-depth research on three text classification tasks: sentiment analysis, offensive language identification, and news topic classification. The datasets used are Stanford Sentiment Treebank (SST-2) 4-1 , Offensive Language Identification Dataset (OLID) 4-2 , and AG’s News 4-3 . Table 2 lists the detailed information of the three datasets.
Table 2: Details of the three evaluation datasets. “Classes” indicates the number of classes and labels. “Avg.#W” represents the average sentence length (number of words). “Train” and “Test” refer to the numbers of samples in the training set and test set, respectively.
@@ 4.1.1 Evaluation Datasets
4.1.2 Victim Model
In recent years, pre-trained large language models (PLMs), represented by GPT-4 gpt4 , have made breakthrough developments. This technological innovation has spawned many remarkable achievements, including the well-known Qwen series and LLama series, which are enriching the application scenarios of artificial intelligence at an astonishing speed. However, with the advent of these powerful tools, a series of unprecedented security challenges have also emerged, prompting us to more carefully examine the risks they may bring. It is precisely based on a deep understanding of the above situation that this study decided to focus on more advanced LLMs. In the end, we determined that the three victim models were Qwen2-72B-Instruct 4-4 , LLama3-8B-Instruct llama3 , and LLama3.2-3B-Instruct llama3 , as shown in Table 3. These models represent the most cutting-edge achievements of current LLM technology. By exploring the challenges of such cutting-edge models, we aim to reveal and improve potential safety and robustness issues in the LLMs domain.
Table 3: Details of the three victim models. “Params” indicates the parameter size of the model, in billions of parameters. “Architecture” refers to the basic architecture type used by the model. “Context length” represents the context length that the model can effectively process, also known as the maximum sequence length. “Token count” indicates the number of tokens in the pre-training data (where “T” means “trillion”). “Release Date” shows the exact date when the model was released.
@@ 4.1.2 Victim Model
4.1.3 Baseline Methods
We select three representative textual backdoor attack methods as baseline methods. (1) BadNet Badnets , originally a backdoor attack method for images, was later cleverly adapted by Kurita and his team to a form suitable for textual attacks RIPPLEWeight . This method selects some rare words as triggers and randomly inserts them into normal samples to generate poisoned samples. (2) InsertSent Insertsent , which uses fixed sentences as triggers and randomly inserts them into normal samples to generate poisoned samples. (3) Syntactic Hiddenkiller , which uses syntactic structures as triggers for textual backdoor attacks and conducts extensive experiments to demonstrate that the syntax-based attack method can achieve comparable attack performance to insertion-based methods, but with higher stealth and stronger defense capabilities.
4.1.4 Evaluation Metrics
Building on previous work Insertsent , RIPPLEWeight , Hiddenkiller , we introduce two evaluation metrics to evaluate and compare the effects of backdoor attacks. (1) Clean Accuracy (CACC): This is defined as the accuracy of the model on the original, undisturbed test data without being attacked, poisoned, or contaminated in any way. This is used to evaluate the performance of the backdoor model on clean data. (2) Attack Success Rate (ASR): This is an important metric for measuring the effectiveness of backdoor attacks. Specifically, ASR calculates the ratio of the number of times the model outputs the attacker-specified result when encountering a test sample with an embedded trigger to the total number of attacks.
4.1.5 Implementation details
The experiments in this study were mainly completed in the following hardware and software environments. In terms of GPU, most experiments used two NVIDIA RTX A6000 graphics cards (48GB of video memory each), and a small number of experiments were conducted on a dual-card NVIDIA GeForce RTX 3090 (24GB of video memory each) platform. The CPU uses AMD Ryzen Threadripper PRO 5975WX, with 32 cores and 64 threads, and a main frequency of 3.6 GHz. The memory system is configured with 4 DDR4 memory with a total of 128GB and a frequency of 3200 MHz. The operating system is Ubuntu 22.04 LTS. The model fine-tuning stage is implemented based on the LLaMAFactory framework, which supports efficient adaptation and training optimization of large language models.
According to the conclusions drawn from previous work Insertsent , the choice of target label has almost no impact on the results of backdoor attacks. Therefore, we set the target labels for the poisoned samples of these three datasets as ”Positive”, ”Offense” and ”World” respectively.
@@ 4.1.5 Implementation details
In backdoor training, we use the adamw_torch 5-1 optimizer and set the initial learning rate to 5.0e-4 (we use the most advanced fine-tuning architecture to fine-tune the victim model, and in the process decided not to use the learning rate setting in the baseline method. In previous studies, for example, Hiddenkiller used 2e-5 as the initial learning rate; however, according to our experimental reproduction results, the use of such a learning rate has led to performance regression. Therefore, we choose to explore more efficient learning strategies to achieve the best model performance). The ratio of the learning rate warmup phase is 0.1, and the type of the learning rate scheduler is ”cosine”. The learning rate of each batch is decayed by cosine annealing, The process is formulated by (6) 5-2 . In addition, we train the three models Qwen2-72B-Instruct, LLama3-8B-Instruct, and LLama3.2-3B-Instruct for different rounds. Specifically, they are trained for 5, 4, and 3 rounds respectively. The other hyperparameters and training settings of the baseline are the same as its original implementation.
@@
Table 4: Attack performance of different backdoor attack methods. “Benign” indicates a benign model without a backdoor. Bold numbers indicate significant improvement over the control group, and underlined numbers indicate no significant difference compared to the control group.
@@
4.2 Backdoor Attack Results
Table 4 shows the experimental results of implementing different backdoor attack methods on three victim models based on three different datasets. It can be seen from the data that these attack methods have achieved extremely high ASR on all victim models, while having little impact on the CACC of the model, which reveals the vulnerability of LLMs when facing backdoor attacks. Compared with the other three baseline methods, the dual-trigger attack method using dual triggers significantly outperforms the syntax-based method in terms of attack performance and achieves a comparable attack performance (almost 100% ASR) with the insertion-based method. Specifically, among the three datasets, dual-trigger performs best on AG’s News, while the effect on the SST-2 dataset is relatively weak. According to our analysis, the dataset size may be an important factor affecting the attack performance of dual-trigger. Given that dual-trigger uses abstract syntactic features and subjunctive mood, a larger amount of data may be required to optimize the attack effect during the backdoor training process. Among the three datasets, AG’s News has the largest data size and SST-2 has the smallest data size.
4.3 Results of the attack with only the subjunctive activated
According to the data in Table 4, the dual-trigger backdoor attack method has made significant progress compared to the baseline method Syntactic. However, these experimental results are based on backdoor test datasets embedded with two layers of triggers. In order to more comprehensively evaluate the advantages of the dual-trigger method, we further experimented on test datasets with only a single trigger embedded.
Specifically, we fine-tuned the customized LLM to create a subjunctive-only version for each test dataset, that is, embedding the subjunctive as a trigger into the original sample. The attack results with only a single trigger activated are shown in Table 5. We can see that when faced with test samples with only the subjunctive embedded, the dual-trigger method still achieved a very high ASR. This shows that the dual-trigger backdoor attack method not only achieves a high ASR, but also has two trigger mechanisms that can be independently enabled during the attack phase, which improves the flexibility of the trigger style and the ability to resist common backdoor defense strategies.
Table 5: In the test phase, the ASR of Dual-Trigger is evaluated using test datasets containing only the subjunctive mood trigger
@@ 4.3 Results of the attack with only the subjunctive activated
4.4 Impact of poisoning rate
@@
In order to evaluate the impact of defense measures on the poisoning rate, this paper conducted a special study based on the AG’s News dataset. As shown in Fig. 5(a), for the Qwen2-72B-Instruct model, when the poisoning rate reached 10%, its ASR and CACC performed well; however, after adopting the Onion defense strategy, CACC dropped significantly from 89.91% at point A to 34.93% at point B. In contrast, when the poisoning rate was only 5%, CACC was relatively less affected by Onion defense, so we determined 5% as its final poisoning rate. According to Fig. 5(b) and 5(c), for the other two models, we determined 10% as their final poisoning rate.
@@
5 Attack Invisibility and Robustness against defense strategies
In this section, we will explore in depth the invisibility of different backdoor attack methods and their robustness against defense strategies. Among them, the invisibility of the trigger is a key factor in backdoor attacks, refers to the fact that the poisoned samples are indistinguishable from normal samples, making it difficult to distinguish them 6-1 . Highly invisible design can effectively bypass manual or automated data review processes, greatly reducing the risk of being detected and removed. Given that many defense strategies rely on meticulous data review, there is a close connection between the invisibility of backdoor attacks and their robustness against defense strategies. In order to evaluate the invisibility of dual-trigger, we used several automated data quality evaluation methods to evaluate the differences between the poisoned samples it generated and normal samples.
5.1 Automated Data Quality Evaluation
@@
These three automated data quality evaluation strategies specifically include: semantic similarity analysis (SSA) based on the ”all-MiniLM-L6-v2”444”all-MiniLM-L6-v2” is a pre-trained language model provided by Microsoft, belonging to the MiniLM series. It is a small but efficient transformer model, especially suitable for tasks such as semantic text similarity, text classification, and information retrieval. language model, perplexity (PPL) evaluated using the GPT-2 model 6-2 , and comprehensive text quality checks using the LanguageTool [^5] https://www.languagetool.org. The latter is further divided into four dimensions: overall text score (TS), number of spelling errors (SEN), number of grammatical errors (GEN), and number of style issues (SIN). When using LanguageTool to evaluate text quality, for each backdoor attack method, the number of GEN is calculated based on 600 samples randomly selected from the corresponding poisoned test dataset. The other three indicators are average values calculated based on the overall performance of the entire poisoned test dataset. This comprehensive consideration method aims to provide a more comprehensive and accurate text quality evaluation system.
The results of six types of indicators calculated according to the three automatic data quality evaluation strategies are shown in Table 6. First, in terms of the SSA indicator (its calculation result is a number between 0 and 1. When this number exceeds 0.8, it is generally considered that two sentences have a high degree of semantic consistency), we can see that the insertion-based backdoor attack method achieved a very high score, because the insertion-based method does not make any changes to other components in the sentence except inserting rare words or short sentences. Our method achieved a comparable performance to the insertion-based method on SSA and was much better than the syntax-based method. In addition, among the 15 evaluation results of PPL, TS, etc. on the three datasets, our poisoned dataset achieved the best results in 14 of them. This fully demonstrates that the method we adopted in creating training datasets with embedded backdoors and test datasets after defense processing is very effective.
Table 6: Results of automatic quality evaluation on poisoned datasets for different backdoor attack methods. SSA represents the semantic similarity analysis score; PPL represents the perplexity; TS is the overall score of the text; SEN records the number of spelling errors; GEN reflects the number of grammatical errors; and SIN indicates the number of style issues. For each backdoor attack method, the number of GEN is calculated based on 600 samples randomly selected from the corresponding poisoned test dataset.
@@
5.2 Robustness against defense strategies
Nowadays, people have fully realized the threat posed by backdoor attacks, and have proposed a series of defense strategies to cope with this challenge. At present, a variety of textual backdoor defense strategies have been widely used: (1) ONION ONION : It is based on test sample inspection and can be applied to any victim model. (2) Back-translation Paraphrasing: The core idea of this method is to first translate the data sample into Chinese, and then translate it back to English, hoping that the trigger factors embedded in the test sample can be eliminated through paraphrasing. (3) Syntactic Structure Alteration: This scheme relies on a specific algorithm to process each sample and reconstruct it into a version with an extremely common syntactic structure ”S(NP)(VP)(.)”. The purpose of this is to reduce or even eliminate those special syntactic features that may be used to activate hidden backdoor attacks. In this section, we will analyze and compare the resistance of different backdoor attack methods to the above defense strategies one by one.
5.2.1 Robustness against ONION
Table 7 shows the attack results of different backdoor methods after applying the ONION defense strategy. From the data in the table, it can be seen that the deployment of ONION has little effect on the CACC of the benign model and the backdoor model, but significantly weakens the ASR of the two insertion-based baseline methods (the average ASR of the two attack methods is reduced by more than 47%). However, for the attack method based on syntax and the dual-trigger attack method, the impact of ONION is relatively limited, with the ASR of the former only decreasing by an average of 6.43%, and the latter by only 2.17%. This shows that our dual-trigger backdoor attack is more resistant to existing such defense measures.
Table 7: Attack performance of different backdoor attack methods under ONION defense (the numbers in brackets are the differences compared with the case without defense).